FOURIER TRANSFORM EMISSION SPECTROSCOPY OF SOME NEW BANDS OF ReN

\underline{P. F. BERNATH}, Department of Chemistry, University of Waterloo, Waterloo, Ont., Canada N2L 3G1 and Department of Chemistry, University of York, Heslington, York, UK, YO10 5DD; R. S. RAM, Department of Chemistry, University of Arizona, Tucson, AZ 85721.

The emission spectrum of ReN has been reinvestigated in the visible region using a Fourier transform spectrometer. Two new bands have been located at 22110 and 22224 cm^{−1} and assigned as the 0^{+} - A1 and 0^{−} - A1 transitions. From the rotational analysis of these bands it has been concluded that the current 0^{+} - A1 transition and the [24.7]0^{+} - X0^{+} transition observed previously by Balfour et al. [J. Mol. Spectrosc. 183, 113-118 (1997)] have an upper state in common. This assignment provides T_{00} = 2596 cm^{−1} for the A1 state. It is likely that the A1 and X0^{+} states are two components of the ^{3}Σ^{−} ground state.